lolokrat:

Geometry help, please!!! Determine if the lines are parallel, perpendicular, or neither. #1) y=1/2x+7 y-8=-2(x+9)

lolokrat:

@Hero

Hero:

1. Put the second equation in the form \(y = mx + b\) 2. Test the slopes using \(m_1m_2 = -1\) If their products = -1 then the lines are perpendicular.

lolokrat:

y=1/2x+7 and y-8=-2(x+9) where do i go from here?

Hero:

I just told you above what to do. Follow the steps.

lolokrat:

I'm unsure of what to plug in

dude:

Like Hero said, isolate y \(y-8=-2(x+9)\) First we want to distribute to make this easier, can you do that?

lolokrat:

y-8=-2(x+9) y-8=-2x+18 y=-2x+10 is this right?

Hero:

How come you didn't do that when I asked?

lolokrat:

i didn't know what you meant when you said m-m1

Hero:

$m_1$ and $m_2$ are the slopes from each line.

dude:

Not quite, you have to do \(-2 \times x \) + \(-2\times 9\) \(-2 \times x \) is -2x and \(-2\times 9\) is -18 (You missed the sign transfer) Then you have to add 8 on both sides, right idea \(y-8=-2x-18\) ==> \(y=-2x-10\) Keep the sign and you're good

Hero:

All that matters really are the slopes.

Hero:

You have them now multiply them together. If you get -1 then the lines are perpendicular.

dude:

Well true but I just want to show the whole process, shortening it might be confusing

lolokrat:

I was unsure of the steps, thanks though Hero.

lolokrat:

y=-2x-10 times y=1/2x+7 is y=-2x+-70 ????

dude:

Hero said to multiply the slopes

lolokrat:

y=-1x+-70 ***

lolokrat:

oh

lolokrat:

so it is perpendicular, it =-1

dude:

Yep

dude:

https://www.desmos.com/calculator/mg3bxn27zh Graphically (Not sure if you were allowed to use graphing calculators)

lolokrat:

yeah, can't use graphing calcs

lolokrat:

i have another, can i work it out and you check it?

dude:

Sure

lolokrat:

y+3=5(x-4) y-2=5(x+2) y-2=5(x+2) y+3=5x-20 y=5x+23 -- y-2=5(x+2) y-2=5x+10 y=5x-8 -- 5*5=25 ?

dude:

I'm confused on your work, what did you do

dude:

If I'm reading this right \(y+3=5x-20\\ y=5x+23\) Watch your sign -20-3 ≠ [+]23